# Restless Multi-armed Bandits under Frequency and Window Constraints for Public Service Inspections

### v1

@@ 6.4 Results
In practice, there will be errors in the estimates of the establishments’ transition probabilities. We perform a sensitivity analysis on this error: adding noise to the parameter estimates from a Normal distribution with a mean of 0 and variances ranging from 0.05 to 0.2. We compare our method (Opt, Opt, =1) with the baseline (Opt, IP, =1). The result shows the difference of rewards before and after adding noise in Table 1. We find that zero-mean Gaussian noise in the exactly one inspection setting in the synthetic domain reduces reward by around 100 points for both (Opt, Opt, =1) and (Opt, IP, =1), and we find that the reward is reduced by the same or less in (Opt, Opt, =1) vs. (Opt, IP, =1).
@@ 7 Conclusions
We present an RMAB-based method to solve scheduling problems under real-world city service scheduling constraints. Both synthetic data results and those using real food inspection data from CDPH suggest that our methods for explicitly modeling constraints and optimizing action windows are critical for RMABs to have an impact in this setting. We hope our work paves the way for applying RMABs to other critical infrastructure maintenance and public service problems under constraints.
Appendix A Appendix
A.1 Algorithm Computational Costs
@@
We simultaneously compute Whittle indices for all states of each arm using binary search over subsidies with the tolerance $10^{-6}$. All experiments are run on a single core of AMD EPYC 7643 (Milan) processors (2.3 GHz). For an RMAB with 1000 arms, computing Whittle Indices takes around 1000s and the baseline scheduling IP takes around 100s for one period. Details of the running time of policies can be seen in the supplement. Optimizing policies consumes around an order of magnitude more computational time than the baseline, but we can reuse the Whittle indices for window optimization for scheduling, and as a result, optimizing windows and schedules together requires negligibly more computation time than performing either optimization individually. RAM consumption is low for all policies, less than 500MB.
A.2 Weighted $b$-Matching
@@
The lookahead planning algorithm we develop will be reducible to variants of the weighted $b$-matching problem Schrijver and others (2003). A weighted $b$-matching instance is described by an undirected graph $G=(V,E)$, an edge weight vector $w:E\rightarrow\mathbb{R}$, and a non-negative $b$ vector $b:V\rightarrow\mathbb{N}_{+}$. The objective in a maximum weight $b$-matching is to find a set of edges $x$ with maximum weight, subject to the constraint that only $b(v)$ edges that are adjacent to node $v$ can be selected. Formally,
@@
Weighted $b$-matchings can be solved in polynomial time, e.g., in $O(|V|^{2}\max_{v}b(v))$ Pulleyblank (1973).
A more challenging weighted $b$-matching variant is weighted bipartite $b$-matching Chen et al. (2016). In this variant, graph nodes are partitioned into a right set $U$ and left set $V$, and there are no edges within each partition. Nodes in the left (resp., right) set have maximum matching cardinality $L^{+}$ (resp., $R^{+}$) and minimum cardinality $L^{-}$ (resp., $R^{-}$). Under these constraints, finding a maximum weight $b$-matching is NP-hard.
A.2.1 Proof For Proposition 1

### v2

@@ 6.4 Results
Figure 4: The reward drop caused by introducing surprise inspections at a 1% rate compared in the synthetic domain relative to the rewards in Figure 2.
@@
In practice, there will be errors in the estimates of the establishments’ transition probabilities. We perform a sensitivity analysis on this error via adding Gaussian noises on the parameters. We compare our method (Opt, Opt, =1) with the baseline (Opt, IP, =1). The result shows the difference of rewards before and after adding noise in Table 1. We find that zero-mean Gaussian noise in the exactly one inspection setting in the synthetic domain reduces reward by around 100 points for both (Opt, Opt, =1) and (Opt, IP, =1), and we find that the reward is reduced by the same or less in (Opt, Opt, =1) vs. (Opt, IP, =1).
CDPH is also responsible for unscheduled inspections, e.g., in response to complaints. We conducted simulation experiment assuming each food establishment has 1% probability of an unscheduled inspections and evaluated the impact to the rewards to the synthetic setting of Fig. 2 and show the results in Fig. 4. We saw both our window and scheduling optimization shows robustness compared with IP and Random policies. With =1 constraint, a regular IP would suffer around 2.6% reward losses while ours only sacrifices 1.9%. If =1 is removed, reward loss would drop less than 0.5% and the gap between our method and IP is narrowing, showing that strict frequency constraints have an essential impact to rewards.
@@ 7 Conclusions
We propose an RMAB-based approach for real-world urban service scheduling. Experiments on synthetic data and Chicago Department of Public Health inspection data show that explicitly modeling constraints and optimizing action windows are crucial for RMAB effectiveness. This work paves the way for applying RMABs to other constrained public service and infrastructure maintenance problems.
Appendix A Implementation Details
Additional details and implementation code are available at: https://github.com/Peter-beeler/TimeConstrainedBandit.
Appendix B Inspection Window and MDP Details
We argue that an unanticipated inspection assignment(like ours) is essential in such a public health check. A study by surprise_ins exploits variation in inspection predictability among co-located food-service establishments in LA County to assess how anticipation affects regulatory compliance. Establishments receiving unanticipated sole inspections exhibit significantly worse compliance, with 7.75% more violations and 16.3% more major critical violations than those inspected concurrently with neighboring establishments. This studies show establishments are preparing for inspections meticulously which does not provide insight into critical violations outside the temporal scope of inspections.
To enforce that there are $\eta$ timesteps of no action (sleep) after each action, as in  field_study, requires a factor of $\eta$ more states. We add a counter to the state to record the number of timesteps until the next pull is allowed. When the counter is positive, the effect of an action is the same as the effect of no action.
This means the Whittle heuristic will never select arms outside the action window (or during mandatory sleep) as long as there is some arm with positive action effects. This is because the Whittle index of arms that are not eligible to be pulled is zero as arms outside the action window (or during mandatory sleep) have no advantage for the active action over the passive action. If they are selected anyway, the agent can discard these actions to no ill effect.
Appendix C Algorithm Computational Costs
Policy Time (sec)
(Rdm, IP, =1) $88.49\pm 4.05$
(Opt, IP, =1) $1070.21\pm 44.54$
(Rdm, Opt, =1) $1106.35\pm 52.31$
(Opt, Opt, =1) $1072.24\pm 45.73$
(Rdm, Opt, $\leq$1) $1028.19\pm 51.97$
(Opt, Opt, $\leq$1) $982.34\pm 44.54$
Table 2: The average running time in seconds for a 1000-arm RMAB over 12 steps.
@@ Appendix C Algorithm Computational Costs
We simultaneously compute Whittle indices for all states of each arm using binary search over subsidies with the tolerance $10^{-6}$. All experiments are run on a single core of AMD EPYC 7643 (Milan) processors (2.3 GHz), on Ohio Supercomputer Center (Ohio_Supercomputer_Center1987-dl). For an RMAB with 1000 arms, computing Whittle Indices takes around 1000s and the baseline scheduling IP takes around 100s for one period. Details of the running time of policies can be seen in Table 2. Optimizing policies consumes around an order of magnitude more computational time than the baseline, but we can reuse the Whittle indices for window optimization for scheduling, and as a result, optimizing windows and schedules together requires negligibly more computation time than performing either optimization individually. RAM consumption is low for all policies, less than 500MB.
Appendix D Transition Inferring Model and Data Preparation
Data Preparation
The inputs to the model are a series of features of establishments. We use the same features as previous work on predicting food establishment risks  (foodforecast), which combines data of business licenses, food inspections, crime, garbage complaints, sanitation complaints, weather and sanitation information. The inspection results are shown in the “Violations” column: 0 means no violations and pass, 1 means violations appear and 2 means pass with conditions. In the experiment, both 0 and 2 are merged into a single good state and 1 is the bad state. We group all inspection data by establishments and transfer them to a transition trajectory data which contains original features, start/end states and time spans.
Figure 5: Structure of neural networks of learning transitions of food establishments.
Inferring Transitions
To create a realistic instance, we must infer the transition probabilities from the inspection trajectories for each arm. Because of the small number of inspections per establishment, it is impractical to estimate a transition model for each food establishment. Thus, we used all data to train one single model to learn the transitions for all establishments. A neural network with 2 MLPs is trained to predict the transition matrix $P^{(0)}_{i}$ to maximize the log-likelihood of the data given the transitions. The detailed architecture of the neural network can been seen in Figure 5.
Appendix E Weighted $b$-Matching
@@
The lookahead planning algorithm we develop will be reducible to variants of the weighted $b$-matching problem (b-matching). A weighted $b$-matching instance is described by an undirected graph $G=(V,E)$, an edge weight vector $w:E\rightarrow\mathbb{R}$, and a non-negative $b$ vector $b:V\rightarrow\mathbb{N}_{+}$. The objective in a maximum weight $b$-matching is to find a set of edges $x$ with maximum weight, subject to the constraint that only $b(v)$ edges that are adjacent to node $v$ can be selected. Formally,
@@
Weighted $b$-matchings can be solved in polynomial time, e.g., in $O(|V|^{2}\max_{v}b(v))$ (pulleyblank_matching).
A more challenging weighted $b$-matching variant is weighted bipartite $b$-matching (wbm). In this variant, graph nodes are partitioned into a right set $U$ and left set $V$, and there are no edges within each partition. Nodes in the left (resp., right) set have maximum matching cardinality $L^{+}$ (resp., $R^{+}$) and minimum cardinality $L^{-}$ (resp., $R^{-}$). Under these constraints, finding a maximum weight $b$-matching is NP-hard.
Proof For Proposition 1
@@
Appendix F Additional Experiment Results
F.1 Restore Failure
In our baseline model, we assume perfect restoration following inspection, as the regulatory system prohibits establishments from reopening after critical violations are identified. To assess the robustness of our framework, we examine a scenario in which restoration may fail stochastically. Specifically, we modify the active transition matrix $P_{i}^{(1)}(0,0)$ from 1 to 0.95, representing a 5% probability that an inspection fails to restore an establishment to a compliant state. All other components of the framework, including the Whittle index computation and the optimization window, remain unchanged. We evaluate the policy (Opt, Opt, =1) under both restoration rates. The results show that perfect restoration (100%) yields an reward of 41898.92, while the 95% restoration rate produces 41894.01. This difference of approximately 0.01% is negligible and demonstrates that our algorithm remains robust to small levels of restoration failure.
F.2 Hyper-paprameter Finetuning
[…]
Appendix G Fairness Study
Our work proposes algorithms to optimize inspection allocation and prioritization using historical public food-establishment inspection data. While the primary goal is to improve efficiency and public safety, automated decision systems of this type can also produce equity-related effects, both positive and negative.
Potential benefits
By formalizing inspection scheduling as a transparent, data-driven optimization problem, our framework can help reduce arbitrary or ad-hoc decision making. When paired with publicly available data and interpretable models, this may enhance procedural fairness by making the criteria for public service scheduling more consistent. Moreover, constraints represented by optimization problems ensure the a fair frequency of disturbance for businesses.
Potential risks
However, the approach also inherits the biases present in historical inspection data. Past inspection patterns in cities can reflect structural inequities—for example, in foodforecast’s study, the inspector has a great factor weight in XGBoost result, this may also cause some biases our transition-predicting model though consequences are not supposed to be catastrophic due to our constraints.
Mitigation strategies
To address this, future work should incorporate fairness constraints or equity-aware regularization into the optimization objective, ensuring that inspection frequency or expected penalties are balanced across demographic or geographic subgroups. Periodic auditing and transparency reports could monitor the distribution of inspections and outcomes by neighborhood, establishment size, or ownership demographics. Furthermore, participatory feedback from local stakeholders can guide the definition of fairness criteria that align with community values.
In sum, while our method can improve the effectiveness of public health inspection systems, its deployment must be accompanied by deliberate equity safeguards—including bias audits, and multifaceted evaluations during the operation.
Appendix H Demographic Fairness IP
To study equity and resource-allocation balance, we extend the planning by adding fairness constraints that restrict the allocation of service to each demographic or geographic subgroup (e.g., each neigborhood).
Let $\mathcal{I}$ be the set of agents (establishments), $\mathcal{G}\subseteq\mathcal{I}$ denote a demographic or geographic group (e.g., a ward), and $T$ the planning horizon. For each $i\in\mathcal{I}$ and $t\in\{1,\dots,T\}$. Let $B_{t}$ be the inspection budget at time $t$.
The fairness-constrained look-ahead integer program is
$\max_{a}$ $\sum_{t=1}^{T}\sum_{i\in\mathcal{I}}w_{i,t}\,a_{i,t}$
s.t. $\sum_{i\in\mathcal{I}}a_{i,t}\leq B_{t},$ $\forall t=1,\dots,T,$
$\sum_{i\in\mathcal{G}}\sum_{t=1}^{T}a_{i,t}\geq\lambda\sum_{i\in\mathcal{I}}\sum_{t=1}^{T}a_{i,t},$ $\lambda\in[0,1],$
$a_{i,t}\in\{0,1\},$ $\forall i\in\mathcal{I},\,t=1,\dots,T.$
Constraint (2) enforces that group $\mathcal{G}$ receives at least a $\lambda$-fraction of total inspection attention over the horizon 1:T1\!:\!T, i.e., a demographic-parity or minimum-coverage requirement. For the different fraction requirements in a group, $\mathcal{G}_{k},\;k=1,\dots,K$ with individual targets $\lambda_{k}$, constraint (2) would be:
$\sum_{i\in\mathcal{G}_{k}}\sum_{t=1}^{T}a_{i,t}\geq\lambda_{k}\sum_{i\in\mathcal{I}}\sum_{t=1}^{T}a_{i,t},\qquad\forall k=1,\dots,K.$
Noted our IP look-ahead 4.3 is capable of encoding many such fairness constraints over the look ahead window, illustrating the extensibility of our model.